OpenStudy (anonymous):

use log laws to simplify (1/sqrte)

OpenStudy (anonymous):

u mean\[\log \frac{1}{\sqrt{e}} \]???

Parth (parthkohli):

\[{1 \over \sqrt{e}} \implies {\left(e^{1 \over 2} \right)^{-1}}\]Something like that?

OpenStudy (anonymous):

thanky you again parthkohli another straight answer i better become a fan

Parth (parthkohli):

If it is \(\ln \left| {1 \over \sqrt{e}}\right|\), then it's simplified to,\[\ln \left|\left(e^{1 \over 2} \right)^{-1} \right|\]and it is easy to manage from that point.

OpenStudy (anonymous):

thanks again